OpenStudy (morganmay19):

If the factors of a polynomial are x + 3 and x + 7, what values of x make that polynomial 0? A.-3 and -7 B.1 and 3 C.3 and 7 D.

OpenStudy (ikram002p):

so u have f(x)=(x+3)(x+7)

OpenStudy (ikram002p):

let (x+3)(x+7)=0 then u have two cases case 1:- x+3=0 case 2:- x+7 =0 ok ?

OpenStudy (morganmay19):

thank u

OpenStudy (ikram002p):

np :D